# Deep Learning Innovations for Underwater Waste Detection: An In-Depth Analysis

### v1

@@ 2.1 Challenges in underwater trash detection
The detection and management of marine debris are critical to preserving marine ecosystems. The primary challenge lies in the debris’ small size relative to the vast ocean, its submerged nature, and the potential for concealment on the sea bottom [6]. A comprehensive review has been done on the various existing literature in this domain explaining the problems faced in harsh marine environments, shown in Figure 1
@@ 2.3 Deep Learning’s Pivotal Role in Underwater Debris Detection
Deep learning not only facilitates the detection of marine debris but also enhances the quality of underwater imagery. Notable studies employing deep learning for debris detection include the works of ocean pollution detection[21], identification of floating plastic[22], deep sea debris detection[23, 24], submerged debris detection[25], haze removal[26], localization of trash[27], and custom networks for detection[28]. These efforts demonstrate the efficacy of deep neural networks in identifying underwater debris, with innovations such as the ’YOLOTrashCan’ network emerging as a potent tool for marine debris identification [27, 28].
@@ 2.7 Broadening the Scope of Marine Debris Detection Models
The broad scope of marine debris detection models is evident in the recent contributions to the field, with studies [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61] emphasizing diverse approaches. Guo et al. [36, 37] explore unsupervised underwater image enhancement using transformers, significantly improving image clarity for better detection. A comprehensive review of image analysis methods for microorganism counting, including deep learning approaches, which are essential for understanding the impact of marine debris on microscopic life was performed by Li et al. [38]. Another comparative study on different learning algorithms to classify mangrove species using UAV multispectral images was performed by [39], aiding in the identification of marine debris in coastal ecosystems. Marin et al. [40] investigate deep-feature-based approaches to marine debris classification, revealing the potential of machine learning in accurately categorizing various debris types. Other researchers highlight the importance of accurate classification through a review and case study of marine microdebris[41] and also discuss the use of both remote and in situ devices for assessing marine contaminants[42], emphasizing the role of advanced technology in environmental monitoring.
@@ 3.2 Dataset Comparison
The used dataset [64] was constructed using three major steps: data collection, manual annotation, and data pre-processing by Horizontal flips, rotation, random crops, expanding, and color jittering [54, 67]. This augmented approach to data has demonstrated notable enhancements in detection accuracy. The final dataset contained 10,000 images of size 416x416 mapped into three classes: (i) Underwater Trash, (ii) Rover, and (iii) Biological life. We compare the different existing datasets relevant to this use case based on various factors, such as the number of classes, total images, environmental conditions, and the different classes, from Table 1.
@@ 5.1 Evaluating results
Figure 3: Precision and Recall values for the suggested algorithms YOLOv8n and YOLOv8x with colors Purple and Yellow respectively
Figure 4: Loss values for YOLO v5s,v7,v8n,v8x : Green, Red, Violet, Yellow

### v2

@@ 2.1 Challenges in underwater trash detection
The detection and management of marine debris are critical to preserving marine ecosystems. The primary challenge lies in the debris’ small size relative to the vast ocean, its submerged nature, and the potential for concealment on the sea bottom [6]. Another challenge is the the potential risk to aquatic wildlife. A comprehensive review has been done on the various existing literature in this domain explaining the problems faced in harsh marine environments, shown in Figure 1
@@ 2.3 Deep Learning’s Pivotal Role in Underwater Debris Detection
Deep learning not only facilitates the detection of marine debris but also enhances the quality of underwater imagery. Notable studies employing deep learning for debris detection include the works of ocean pollution detection[21], identification of floating plastic[22], deep sea debris detection[23, 24], submerged debris detection[25], haze removal[26], localization of trash[27], as well as a hybrid loss function considering illumination and detection problems [29] and custom networks for detection[28]. These efforts demonstrate the efficacy of deep neural networks in identifying underwater debris, with innovations such as the ’YOLOTrashCan’ network emerging as a potent tool for marine debris identification [27, 28].
@@ 2.7 Broadening the Scope of Marine Debris Detection Models
The broad scope of marine debris detection models is evident in the recent contributions to the field, with studies emphasizing diverse approaches. Guo et al. [37] explore unsupervised underwater image enhancement using transformers, significantly improving image clarity for better detection. A comprehensive review of image analysis methods for microorganism counting, including deep learning approaches, which are essential for understanding the impact of marine debris on microscopic life was performed by Li et al. [38]. Another comparative study on different learning algorithms to classify mangrove species using UAV multispectral images was performed by [39], aiding in the identification of marine debris in coastal ecosystems. Marin et al. [40] investigate deep-feature-based approaches to marine debris classification, revealing the potential of machine learning in accurately categorizing various debris types. Other researchers highlight the importance of accurate classification through a review and case study of marine microdebris[41] and also discuss the use of both remote and in situ devices for assessing marine contaminants[42], emphasizing the role of advanced technology in environmental monitoring.
@@ 3.2 Dataset Comparison
The used dataset [52] was constructed using three major steps: data collection, manual annotation, and data pre-processing using rotation, random cropping, horizontal flips, expansion, and color jittering [106, 55]. This augmented approach to data has demonstrated notable enhancements in detection accuracy. The final dataset contained 10,000 images of size 416x416 mapped into three classes: (i) Underwater Trash, (ii) Rover, and (iii) Biological life. We compare the different existing datasets relevant to this use case based on various factors, such as the number of classes, total images, environmental conditions, and the different classes, from Table 1.
@@ 5.1 Evaluating results
Figure 3: Precision and Recall values (Y-axis) over 140 epochs (X-axis) for YOLOv8n and YOLOv8x with colors Purple and Yellow respectively
Figure 4: Loss values (Y-axis) over 140 epochs (X-axis) for YOLO v5s,v7,v8n,v8x : Green, Red, Violet and Yellow